Why a -2x =0 ? Where x= infinity . Please explain the cause.

Why a-2x =0 ? Where x= infinity . Please explain the cause.

Grade:12th pass

2 Answers

yash khuran
5 years ago
A^-2x =0 
We know that infinity =1/0 
A^-2(1/0)=A^-2/0=0  
A= 0 ^(1/-2)=0 ( because we can see that ×^2 =7 
Then x=7^1/2 .. and similar it is )
Arun
5 years ago
Dear student
 
This is true because
 
it can be written as 1/a^2x
Now a^2x tends to infinity.
 
hence 1/a^2x will tends to zero.